# Observe Then Act: Asynchronous Active Vision-Action Model for Robotic Manipulation

### v2

@@ IV-C Viewpoint-aware Demo Augmentation
To provide early guidance and reduce exploration, we manually alternate camera and gripper movements to generate a small set of demonstration trajectories. Each demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$ consists of observation-action pairs collected at a fixed rate. To further leverage limited demo data, existing augmentation methods [1, 2, 28] first select key frame indices from the demo trajectory, construct transition samples from the key observation-action pairs, and then resample between key frames to generate additional transitions. However, they cannot handle the observation variance caused by viewpoint shifts. We define the key frame discovery function as $KF:\tau_{\text{d}}\rightarrow\big{\{}(k_{j}^{\text{c}},k_{j}^{\text{g}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$, identifying key frames for both viewpoint and gripper movements. Gripper key frames $k_{j}^{\text{g}}$ are marked when the gripper is stationary or its closure state changes, while key viewpoints are defined as the first stable viewpoint between two gripper key frames, indexed by $k_{j}^{\text{c}}$. The key frame indices satisfy $10\leq k_{j-1}^{\text{c}}\leq k_{j-1}^{\text{g}}\leq k_{j}^{\text{c}}\leq k_{j}^{\text{g}}\leq N-$ 1. As shown in Fig.5, raw transition samples $\text{T}_{j}$ are then constructed as formalized in Eq.(5).
$\begin{split}\text{T}_{j}&=\{\text{T}_{j}^{\text{v}},\text{T}_{j}^{\text{p}},r_{j}\}\\ &=\{(\mathbf{o}_{k_{j-1}^{\text{g}}},\mathbf{a}_{k_{j-1}^{\text{g}}}^{\text{c}},\mathbf{o}_{k_{j}^{\text{c}}}),(\mathbf{o}_{k_{j}^{\text{c}}},\mathbf{a}_{k_{j}^{\text{c}}}^{\text{g}},\mathbf{o}_{k_{j}^{\text{g}}}),r_{j}\}\end{split}$ (5)
To generate additional augmented transitions, we uniformly sample within the intervals $k_{j-1}^{\text{g}}\sim k_{j}^{\text{c}}$ and $k_{j}^{\text{c}}\sim k_{j}^{\text{g}}$, obtaining the augmented indices$(k_{j-1}^{\text{g}})^{\text{aug}}$ and $(k_{j}^{\text{c}})^{\text{aug}}$. The resulting augmented transition $\text{T}_{j}^{\text{aug}}$ is formalized in Eq.(6).
$\begin{split}\text{T}_{j}^{\text{aug}}&=\{(\text{T}_{j}^{\text{v}})^{\text{aug}},(\text{T}_{j}^{\text{p}})^{\text{aug}},r_{j}\}\\ \end{split}$ (6)
Both $\text{T}_{j}$ and $\text{T}_{j}^{\text{aug}}$ are ultimately stored in the replay buffer as early training experiences for the model.
Figure 5: Demo trajectory augmentation.For each demo trajectory, we uniformly sample observations(represented by hatched circles) between both viewpoint key frames(solid blue circles) and gripper key frames(solid green circles) to generate additional transition samples, enabling viewpoint-aware demonstration augmentation.
@@ IV-D Task-agnostic Auxiliary Reward
We use sparse reward to avoid designing task-specific rewards. However, since sparse reward only provides feedback at the end of an episode, it hinders efficient sensor-motor learning. Here we introduce two task-agnostic auxiliary rewards to provide real-time guidance to the training process.
To encourage interaction with the environment and accurate environment modeling, we propose the Interaction Encouragement Reward, denoted as $r_{\text{i}}$ As illustrated in Fig.6, $r_{\text{i}}$ provides positive signals when the NBV policy selects a reachable, non-empty ROI and penalizes actions targeting empty or unreachable areas due to execution failures. The specific value assignment for $r_{\text{i}}$ is detailed in Eq.(7).
$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI unreachable}\end{cases}$ (7)
To encourage the model to select viewpoints that capture more structural information from the ROI, we introduce the Entropy Reduction Reward($r_{\text{e}}$), inspired by [21]. The goal is to reduce occupancy entropy within the ROI after viewpoint adjustment, as shown in Fig.6. Assuming the ROI contains $N$ voxels, the occupancy entropy of the $i$-th voxel($e_{i}$) is calculated using the Shannon entropy function(Eq.(8)), where $p_{\text{o}}(i)$ represents its occupancy probability. If a voxel is occluded, $p_{\text{o}}$ is set to 0.5.
$e_{i}=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-\big{(}1-p_{\text{o}}(i)\big{)}\log_{2}\big{(}1-p_{\text{o}}(i)\big{)}$ (8)
The total occupancy entropy change of the ROI from time $t$ to $t^{\prime}$ is expressed as:
$\Delta E=\frac{1}{N}\sum_{i=1}^{N}e^{t}_{i}-\frac{1}{N}\sum_{i=1}^{N}e^{t^{\prime}}_{i}=E_{t}-E_{t^{\prime}}$ (9)
We define reward $r_{\text{e}}=k\cdot\Delta E$, where $k$ is empirically set to 1.0. The auxiliary reward $r_{\text{i}}$ requires cooperation between the NBV and NBP policies, while $r_{\text{e}}$ is exclusive to the NBV policy. Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, and the NBV reward is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion. These task-agnostic rewards generalize across tasks without the need for manual tuning.
Figure 6: Viewpoint adjustment and ROI selection in a 2D simplified schematic scene. The model is encouraged to select non-empty ROIs and reduce occluded voxels from the adjusted viewpoint $\mathbf{v}_{t^{\prime}}$.
@@ V Experiment Setup
Task Setting. We trained and evaluated our model on RLBench[9], a widely-used public benchmark for robotic manipulation in reinforcement learning. Task selection followed two criteria: 1) each task must include at least one viewpoint with significant visual occlusion, and 2) no single fixed or wrist-mounted camera provides consistently low occlusion across selected tasks. As the ROI shifts throughout the manipulation process, and there may be phases without distinct objects(e.g., midair gripper movements), defining occlusion rates based on object segmentation, as done in previous methods[14, 21], is impractical. To address this, we define Task Occlusion Rate(TOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur, based on observations from that viewpoint across $L=100$ demo trajectories, as described in Eq.(10). Here, $E^{l}_{t^{\prime}_{k}}$ represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$ after viewpoint adjustment where the ROI center is set to the gripper’s goal position in the demo observation. To compare the occlusion levels across selected tasks, we assume that the human-demoed active viewpoint contains the strongest task-specific prior knowledge. Thus, we can obtain the Relative TOR by dividing the TOR of the selected viewpoint by the TOR of the human-demoed active viewpoint(oracle).
$\text{TOR}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}E^{l}_{t_{k}^{\prime}}$ (10)
Fig.7 presents the relative TOR statistics for 8 selected RLbench tasks under overhead, front, wrist-mounted cameras as well as the human-demoed active camera.
Figure 7: Reletive task occlusion rate of selected tasks. We use the active viewpoint occlusion rate of 100 demonstration trajectories for each task as a benchmark.
@@
Baselines. We compared our model against two baseline categories: 1) visual manipulation model that utilize static third-person and wrist-mounted cameras, and 2) active third-person camera-based models. For the first category, we compared our model against C2FARM [2] with static front, overhead and wrist-mounted camera settings to explore the potential effectiveness of our active vision-action model in handling tasks with limited observability. For the second category, we compared our model with Behavior Cloning(BC) to test whether it can learn new sensor-motor skills beyond demonstration imitation. We also tested a Random viewpoint model [3], which shares the same architecture as ours but with the NBV policy network disabled(learning rate set to zero). This setup allows us to assess how much performance improvement is achieved through active viewpoint adjustment as opposed to random choices.
@@
Training Setting. We collected 30 demo trajectories for each task and used them as prior knowledge for training. C2FARM was implemented based on its official release and trained with sparse task rewards. The Behavior Cloning model was trained with MSE loss, and the Random model used the same reward structure as our model(SectionIV-D). All models were trained using Adam with a learning rate of $0.0005$ for 40,000 updates, with a batch size of 128. Experiments were conducted on a machine with 4 NVIDIA RTX 4090 GPUs.
@@
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
C2FARM(wrist) 0.93±0.02 0.01±0.01 0.11±0.02 0.44±0.07 0.90±0.02 0.00±0.00 0.52±0.01 0.30±0.02 0.40±0.35
C2FARM(overhead) 0.87±0.02 0.00±0.00 0.28±0.05 0.03±0.01 0.89±0.03 0.61±0.04 0.24±0.03 0.89±0.04 0.48±0.36
C2FARM(front) 0.26±0.03 0.28±0.04 0.68±0.04 0.37±0.04 0.86±0.06 0.91±0.02 0.66±0.06 0.00±0.00 0.50±0.30
@@
Models Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
C2FARM(wrist) 3.73±0.18 9.96±0.04 9.27±0.20 7.06±0.49 3.64±0.23 9.98±0.03 6.39±0.11 7.75±0.17 7.22±2.39
C2FARM(overhead) 5.21±0.17 10.00±0.01 8.50±0.28 9.78±0.07 4.13±0.23 6.52±0.21 8.46±0.21 3.93±0.22 7.07±2.30
C2FARM(front) 8.43±0.15 8.66±0.18 5.91±0.24 8.04±0.20 4.49±0.40 3.22±0.13 5.63±0.41 9.99±0.02 6.80±2.20
@@
VI Result
Main Performance. We first analyzed the average task success rate(SR) and episode length(EL) across different models to evaluate overall sensor-motor performance. Table I shows that our model achieved the highest success rate, outperforming the second-best by 21%, especially in occluded tasks like meat_on_grill, meat_off_grill and unplug_charger. Similarly, Table II shows that our model had the shortest average manipulation steps, outperforming the second-best by 28.2% and showing the least variance in step counts.
Fixed and wrist-mounted cameras performed well in low-occlusion tasks, such as basketball_in_hoop, press_switch, put_rubbish_in_bin, C2FARM occasionally matching or surpassing our model. However, their performance dropped sharply in occluded scenarios, leading to a higher variance. In contrast, our model remained robust in occluded tasks. Although the BC model demonstrated some ability to adjust its viewpoint(as detailed below), it struggled with limited training samples, achieving only a 14% success rate. The random viewpoint model performed the worst across all tasks, highlighting its inability to learn sensor-motor coordination.
Figure 8: Partial rollout cases and viewpoint distribution. (a) The image observations corresponding to the first three inferred viewpoints $\mathbf{v}^{*}$ and the gripper actions $\mathbf{g}^{*}$ for four representative tasks during the test episodes. (b) Red circles indicate manually set viewpoints from 30 demo trajectories, while blue triangles represent the learned viewpoint positions from successful episodes across 100 test runs.
Active Perception Strategy Analysis. We further analyzed the perception strategies of all models using two metrics: Average Single Step Information Gain(ASSIG) and Average Initial Information Gain(AIIG). ASSIG captures the average reduction in ROI entropy before and after a single viewpoint shift, while AIIG measures the entropy reduction compared to the initial viewpoint.
$\text{ASSIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}(E^{l}_{t_{k}}-E^{l}_{t^{\prime}_{k}})$ (11)
$\text{AIIG}=\frac{1}{L}\sum_{l=1}^{L}\sum_{k=0}^{K_{l}}([E^{l}_{t_{k}}]_{\mathbf{v}_{t_{0}}}-E^{l}_{t^{\prime}_{k}})$ (12)
Where $L=100$ denotes the number of testing episodes, $K_{l}$ represents the length of episode $l$, and $E^{l}_{t_{k}}$ is defined in Eq.(9). The term $[E^{l}_{t_{k}}]_{\mathbf{v}{t_{0}}}$ represents the visual entropy of the current ROI from the episode’s initial viewpoint. As shown in Table III, our model significantly enhances ROI observation by adjusting viewpoints, outperforming the BC model by 2.8% in ASSIG and 1.2% in AIIG. In contrast, static front and overhead cameras, which cannot shift viewpoints, fail to achieve information gain. The wrist-mounted camera, unable to move independently of the gripper, shows zero ASSIG and is unsuitable for AIIG computation.
Fig.8(b) compares the distribution of viewpoints learned by our model with the manually set active viewpoints from the demo trajectories used during training. Unlike demo patterns, our model’s learned viewpoints are more concentrated in some tasks(Basketball_in_hoop, unplug_charger) and more varied in others(meat_on_grill, open_drawer), indicating it learned to select task-relevant viewpoints rather than simply imitating demo data. The camera observations at key moments(Fig. 8) further showcase our model’s active visual perception capabilities.
Figure 9: Proportion of success, timeouts, and execution failures across test episodes for different models.
@@ VI Result
TABLE III: Comparison of information gain and interaction capacity across different models.
Models ASSIG $\uparrow$ (avg/max) AIIG $\uparrow$ (avg/max) First $\downarrow$ Interaction Non- $\downarrow$ interaction
C2FARM(wrist) 0.0/0.0 N/A 3.21±1.82 4.86±2.70
C2FARM(overhead) 0.0/0.0 0.0/0.0 2.97±1.70 4.41±2.77
C2FARM(front) 0.0/0.0 0.0/0.0 3.41±2.28 4.17±2.74
@@
Environmental Modeling and Interactive Capabilities. We evaluated the models’ environment modeling and interaction capabilities by recording success, timeout(TO), and execution failure(EF) rates over 100 test episodes(Fig.9). The C2FARM model exhibited a high failure rate(up to 42%) due to trajectory execution failures. Similarly, the random viewpoint model saw the majority of its test episodes end in failure because of execution errors. Although only 4% of the BC model’s episodes ended due to execution errors, 82% failed due to timeouts. In contrast, our model showed a 13% failure rate from trajectory errors and only 9% from timeouts, indicating effective integration of low-level motion execution.
As shown in Table III, we also analyzed the average steps to first interaction(FI) and non-interactive steps(NI) per model. The BC model exhibited frequent non-interactive steps, correlating with its high timeout rate, suggesting difficulty learning precise manipulations from limited demonstrations. In contrast, our model initiated interactions more quickly, reducing non-interactive steps, and demonstrating superior accuracy in modeling environmental dynamics and structural understanding after training.
Ablation Study. We evaluated the impact of three strategies on manipulation performance:(1) viewpoint-centric voxel alignment(Section IV-B), (2) viewpoint-aware demo augmentation(Section IV-C), and (3) task-agnostic auxiliary rewards(Section IV-D). Separate ablation tests were conducted for each, as well as a combined test removing all strategies.
Table IV shows that each strategy effectively improved success rates(SR), with voxel alignment and viewpoint-aware augmentation contributing the most, as success rates dropped by approximately 40% without them and auxiliary rewards added around 6% improvement.
Beyond success rate improvements, all three strategies contributed to reductions in episode length(EL), timeouts rate(TO), execution failures(EF), and the average number of non-interactive steps(NI) per episode. This demonstrates their broader importance to the model’s effectiveness. Interestingly, while the voxel alignment strategy significantly boosts overall performance, we found it leads to a reduction in ASSIG score. We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations. However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness.
Ablating all strategies led to a sharp decline, reducing the average success rate to 22%, underscoring the synergistic effect of combining these strategies for optimal performance.

### v3

@@ IV-C Viewpoint-aware Demo Augmentation
During demonstration, the viewpoint and gripper movements are manually alternated asynchronously, staying stationary while the other is moving. These paired stationary viewpoints and gripper positions are called key actions, corresponding to the actions that need to be inferred by the NBV and NBP policies. The paired data frame index of these key actions in one demo trajectory $\tau_{\text{d}}=\{(\mathbf{o}_{i},\mathbf{a}_{i})\}^{N}$, are denoted as $\big{\{}(k^{\text{v}},k^{\text{p}})\big{\}}^{M}\subset\mathbb{R}_{+}^{2}$ Naively, we can use these key observation-action pairs to construct a prior transition sample set $\{\text{T}\}^{M}$ for active sensor-motor training, where T defined as
$\text{T}=\{\text{T}^{\text{v}},\text{T}^{\text{p}},r\}=\{(\mathbf{o}_{k^{\text{init}}},\mathbf{a}_{k^{\text{v}}},\mathbf{o}_{k^{\text{v}}}),(\mathbf{o}_{k^{\text{v}}},\mathbf{a}_{k^{\text{p}}},\mathbf{o}_{k^{\text{p}}}),r\}.$ (5)
However, the number of these raw transition samples is quite limited due to the small demonstration set. To fully utilize the demo trajectory data, we uniformly re-sample the initial observations of both $\text{T}^{\text{v}}$ and $\text{T}^{\text{p}}$ to construct additional transition samples for policy training, as shown in Fig.5. The resulting additional transitions is formalized as
$\text{T}_{\text{a}}=\{\text{T}^{\text{v}}_{\text{a}},\text{T}^{\text{p}}_{\text{a}},r\}=\{(\mathbf{o}_{k^{\text{init}}_{\text{a}}},\mathbf{a}_{k^{\text{v}}},\mathbf{o}_{k^{\text{v}}_{\text{a}}}),(\mathbf{o}_{k^{\text{v}}_{\text{a}}},\mathbf{a}_{k^{\text{p}}},\mathbf{o}_{k^{\text{p}}}),r\}.$ (6)
Both T and $\text{T}_{\text{a}}$ are ultimately stored in the replay buffer as early training experiences for the model.
@@ IV-D Task-agnostic Auxiliary Reward
Relying solely on a centralized sparse reward to coordinate the collaboration between the NBV and NBP agents may result in only one agent being active while the other remains “lazy”, leading to a non-optimal equilibrium between the two agents [36]. To mitigate these issues and provide real-time guidance during the training process, we introduce two task-agnostic auxiliary rewards.
We introduce the Interaction Encouragement Reward $r_{\text{i}}$ to encourage the robot to interact with non-empty ROI areas containing scene entities. If the NBP interaction does not fail due to trajectory planning errors or collisions, and the gripper’s finger center point (GCP) is within the ROI after execution, we refer to this ROI as “reachable”. The specific value assignment for $r_{\text{i}}$ is detailed as
$r_{\text{i}}=\begin{cases}0.02,&\text{ROI reachable and non-empty}\\ 0.0,&\text{ROI reachable but empty}\\ -0.02,&\text{ROI unreachable}.\end{cases}$ (7)
Figure 5: Demo trajectory augmentation. The raw transition T only contains observations and actions at key moments, while the additional augmentative transitions $\text{T}_{\text{a}}$ are constructed from intermediate observations and actions.
We introduce the Entropy Reduction Reward $r_{\text{e}}$ to encourage NBV agent to select viewpoints that capture more determined visual information of the ROI. The reward $r_{\text{e}}$ is defined as $r_{\text{e}}=k\cdot\Delta E$, where $k$ is empirically set to 1.0 and $\Delta E$ is the total occupancy entropy change of the $N$ voxels inside the ROI after viewpoint adjustment. The occupancy entropy of the $i$-th voxel, $e_{i}$, is calculated using the Shannon entropy function [22], shown in Eq.(8), where $p_{\text{o}}(i)$ represents its occupancy probability, and $q_{\text{o}}(i)=1-p_{\text{o}}(i)$.
$\begin{split}\Delta E&=E_{t}-E_{t^{\prime}}=\frac{1}{N}\big{(}\textstyle\sum_{i=1}^{N}e_{i}^{t}-\textstyle\sum_{i=1}^{N}e_{i}^{t^{\prime}}\big{)}\\ e_{i}&=-p_{\text{o}}(i)\log_{2}\big{(}p_{\text{o}}(i)\big{)}-q_{\text{o}}(i)\log_{2}\big{(}q_{\text{o}}(i)\big{)}\end{split}$ (8)
The reward $r_{\text{i}}$ requires cooperation between the NBV and NBP policies, while $r_{\text{e}}$ is exclusive to the NBV policy. Therefore, the reward for updating the NBP policy is expressed as $r_{\text{nbp}}=r_{\text{task}}+r_{\text{i}}$, and the NBV reward is $r_{\text{nbv}}=r_{\text{task}}+r_{\text{e}}+r_{\text{i}}$, where $r_{\text{task}}\in[0,1]$ is determined by task completion. These rewards are unrelated to specific task objectives, thus generalizing across tasks without the need for manual tuning.
@@ V Experiment Setup
Tasks and Initialization Settings. We trained and evaluated our model on RLBench [12], a widely used public benchmark for robotic manipulation. The basic scene of RLBench is illustrated in Fig.3, featuring several pre-set cameras, including fixed front and overhead views, as well as a wrist-mounted camera attached to the gripper. At the start of each episode, the scene objects are randomly positioned within a legal region of $0.67\text{m}\times 0.91\text{m}$ on the tabletop and randomly rotated around the $z$-axis of the world frame, which is perpendicular to the table. This ensures variation in scene initialization, guaranteeing differences in each training and testing episode. We selected 8 tasks (Table I) from RLBench following two criteria: 1) each task must include at least one RLBench default viewpoint with significant visual occlusion, and 2) no single fixed or wrist-mounted camera provides consistently low occlusion across the selected tasks. To evaluate the occlusion level of tasks and viewpoints, we define the Viewpoint Occlusion Rate (VOR) for a given viewpoint as the average occupancy entropy of the ROI where the gripper’s goal actions occur. This is calculated based on observations from that viewpoint across $L=100$ demonstration trajectories, as described in Eq.(9). Here, $E^{l,k}$, represents the ROI occupancy entropy at the $k$-th keyframe in trajectory $l$, after the viewpoint adjustment. Furthermore, we normalize the VOR of all viewpoints with reference to the VOR of the human-demonstrated viewpoint in the same task. This normalization facilitates intuitive comparisons across tasks and viewpoints. The results are shown in Fig.7.
$\text{VOR}=\textstyle\frac{1}{L}\textstyle\sum_{l=1}^{L}\sum_{k=0}^{K}E^{l,k}$ (9)
Figure 6: Viewpoint adjustment and ROI selection in a 2D schematic scene. We define an ROI as non-empty only if it contains at least one voxel with an occupancy probability of $p_{\text{o}}=1.0$; otherwise, it is considered empty.
@@
Baselines. We compared our model against three baseline categories: 1) visual manipulation models that utilize static third-person and wrist-mounted cameras, 2) active third-person camera-based models, and 3) Oracle baseline, which utilizes fully fused image from all default camera views in RLBench. For the first category, we compared our model to C2FARM (C2F) [2] using static front, overhead, wrist and front + wrist (f+w) camera settings to evaluate the effectiveness of our active vision-action model in handling tasks with limited observability. For the second category, we evaluated our model against Behavior Cloning (BC) to test its ability to learn new sensor-motor skills beyond demonstration imitation. Additionally, we included a Random Viewpoint model [3], which shares the same architecture as ours but with the NBV policy network disabled (learning rate set to zero). This setup assesses the performance gain from active viewpoint adjustment versus random selection. For the Oracle baseline, we disabled the NBV viewpoint selection ability since full camera views were always accessible, but retained NBV’s ROI inference capability.
@@
Training Setting. We collected 30 demonstration trajectories for each task, for both ours and baseline methods. used them as prior knowledge for training. The Behavior Cloning model was trained using MSE loss, while the Random model used the same reward as our model but set the NBV learning rate to zero. All models were trained using Adam with a learning rate of $0.0005$ for 40,000 updates. Experiments were conducted on a machine with 4 NVIDIA RTX 4090 GPUs.
VI Result
Main Performance. Main Performance. We first analyzed the task Success Rate (SR) and Episode Length (EL) across all methods on 8 tasks to evaluate their overall performance. Table I shows that our model achieved the highest average success rate, outperforming the second-best single camera setting (C2F+front) by 26%, especially in occluded tasks like meat_on_grill, meat_off_grill and unplug_charger. Similarly, Table II shows that our model had the shortest average episode length, outperforming the second-best single camera setting (C2F+front) by 28.2% , demonstrating its high proficiency in the tasks.
Figure 7: Normalized viewpoint occlusion rate of selected tasks. We consider human-demonstrated viewpoints to have the highest task prior knowledge and use them as the baseline for normalization.
We found that C2F with single camera setting performed well in low-occlusion tasks, such as basketball_in_hoop, press_switch, put_rubbish_in_bin, and it occasionally matching or surpassing our method. However, their performance dropped sharply in occluded scenarios, leading to a higher performance variance. In contrast, our model remained robust in occluded tasks due to its active perception ability. Leveraging complementary information from different viewpoints, C2F with a dual-camera setup (front + wrist) outperforms a single camera significantly but remains slightly below our single active camera setup, with both delivering comparable results. The Oracle baseline, benefiting from global scene perception, achieved the highest success rate and shortest episode length. In addition, although the BC model demonstrated some ability to adjust its viewpoint (as detailed below), it struggled with limited training samples, achieving only a 14% success rate. Meanwhile the random viewpoint policy performed the worst across all tasks, highlighting its inability to learn sensor-motor coordination.
@@ VI Result
Methods Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
C2F (wrist) 0.93±0.02 0.01±0.01 0.11±0.02 0.44±0.07 0.90±0.02 0.00±0.00 0.52±0.01 0.30±0.02 0.40±0.35
C2F (overhead) 0.87±0.02 0.00±0.00 0.28±0.05 0.03±0.01 0.89±0.03 0.61±0.04 0.24±0.03 0.89±0.04 0.48±0.36
C2F (front) 0.26±0.03 0.28±0.04 0.68±0.04 0.37±0.04 0.86±0.06 0.91±0.02 0.66±0.06 0.00±0.00 0.50±0.30
C2F (f+w) 0.95±0.03 0.32±0.04 0.65±0.04 0.55±0.06 0.84±0.04 0.92±0.02 0.85±0.04 0.73±0.01 0.72±0.23
@@
Methods Basketball in Hoop Meat off Grill Meat on Grill Open Drawer Press Switch Put Rubbish in Bin Take Money out Safe Unplug Charger On Average
C2F (wrist) 3.73±0.18 9.96±0.04 9.27±0.20 7.06±0.49 3.64±0.23 9.98±0.03 6.39±0.11 7.75±0.17 7.22±2.39
C2F (overhead) 5.21±0.17 10.00±0.01 8.50±0.28 9.78±0.07 4.13±0.23 6.52±0.21 8.46±0.21 3.93±0.22 7.07±2.30
C2F (front) 8.43±0.15 8.66±0.18 5.91±0.24 8.04±0.20 4.49±0.40 3.22±0.13 5.63±0.41 9.99±0.02 6.80±2.20
C2F (f+w) 3.42±0.21 7.51±0.26 6.14±0.25 6.29±0.35 4.81±0.34 3.13±0.14 5.60±0.17 5.22±0.11 5.27±1.48
@@
Oracle (full views) 3.41±0.13 4.56±0.16 4.33±0.38 4.32±0.32 3.49±0.24 3.08±0.01 3.78±0.04 3.12±0.29 3.76±0.58
Figure 8: Result proportion and partial viewpoint distribution. (a) Proportion of success, timeouts, and execution failures across test episodes for different models. (b)Red squares indicate manually set viewpoints from 30 demo trajectories, while green circles and blue triangles represent the viewpoints learned by the behavior cloning and our model, respectively, from successful episodes across 100 tests.
Active Perception Strategy Analysis. Active perception involves accurately identifying effective ROI regions in the scene and selecting optimal viewpoints to observe them, thereby supporting task completion. We propose Average Single Step Information Gain (ASSIG) and Average Initial Information Gain (AIIG) to quantitatively evaluate the model’s ability to enhance ROI observation quality through viewpoint adjustments. ASSIG measures single-step information gain, while AIIG quantifies improvement relative to the initial viewpoint. They are defined as follows:
$\text{ASSIG}=\textstyle\frac{1}{L}\textstyle\sum_{l=1}^{L}\sum_{t=0}^{T}\Delta E^{l,t}$ (10)
$\text{AIIG}=\textstyle\frac{1}{L}\textstyle\sum_{l=1}^{L}\sum_{t=0}^{T}\Delta E^{l,t}_{\text{init}},$ (11)
Where $L=100$ testing episodes, $T$ is the episode length and $\Delta E$ are defined in Eq.(8). Fig.8(b) shows that both our model and the BC model demonstrate some learned viewpoint adjustment ability. Table.III further confirms this, with our model achieving up to 4.6% and 7.0% higher ASSIG and AIIG scores than the BC model. However, as Fig.8(a) shows, the BC model has significantly more failures due to timeouts, indicating difficulty in accurately localizing ROI regions from limited demos. Additionally, our model learns more task-relevant viewpoints: some tasks (e.g., basketball_in_hoop, put_rubbish_in_bin) show more focused distributions, while others (e.g., meat_on_grill, open_drawer) exhibit greater generalization, reflecting further skill acquisition through interaction rather than mere imitation as seen with BC.
Environmental Modeling and Interactive Capabilities. We evaluated the models’ environment modeling and interaction capabilities by recording Success, Timeout (TO), and Execution Failure (EF) rates over 100 test episodes (Fig.8). The C2F model exhibited a high failure rate (up to 42%) due to trajectory execution failures. Similarly, the Random model saw the majority of its test episodes end in failure because of execution errors. Although only 4% of the BC model’s episodes ended due to execution errors, 82% failed due to timeouts. In contrast, our model showed a 13% failure rate from trajectory errors and only 9% from timeouts, indicating effective integration of low-level motion execution.
As shown in Table III, we also analyzed the average steps to First Interaction (FI) and Non-interactive steps (NI) per model. The BC model exhibited frequent non-interactive steps, correlating with its high timeout rate, suggesting difficulty learning precise ROI location and action skills from limited demonstrations. In contrast, our model initiated interactions more quickly, reducing non-interactive steps, and demonstrating superior accuracy in modeling environmental dynamics and structural understanding after training.
@@
TABLE III: Comparison of information gain and interaction capacity across different methods.
Methods ASSIG $\uparrow$ (avg/max) AIIG $\uparrow$ (avg/max) First $\downarrow$ Interaction Non- $\downarrow$ interaction
C2F (wrist) 0.0/0.0 N/A 3.21±1.82 4.86±2.70
C2F (overhead) 0.0/0.0 0.0/0.0 2.97±1.70 4.41±2.77
C2F (front) 0.0/0.0 0.0/0.0 3.41±2.28 4.17±2.74
C2F (f+w) 0.0/0.0 N/A 1.91±0.55 2.63±1.32
@@
Oracle (full views) 0.0%/0.0% N/A 1.58±1.36 1.86±1.56
Ablation Study. We first evaluated the impact of three strategies on manipulation performance: viewpoint-centric voxel alignment (Sec IV.B), viewpoint-aware demo augmentation (Sec IV.C), and task-agnostic auxiliary rewards (Sec IV.D). Table IV shows that each strategy effectively improved success rates (SR), with voxel alignment and viewpoint-aware augmentation contributing the most, as success rates dropped by approximately 40% without them and auxiliary rewards added around 6% improvement.
Beyond success rate improvements, all three strategies contributed to reductions in Episode Length (EL), Timeouts rate (TO), Execution Failures (EF), and the average number of Non-interactive steps (NI) per episode. This demonstrates their broader importance to the model’s effectiveness. Interestingly, while the voxel alignment strategy significantly boosts overall performance, we found it leads to a reduction in ASSIG score. We hypothesize that without viewpoint alignment, the NBV agent avoids potential issues with out-of-bounds actions and clipping, which may simplify the learning of viewpoint transformations. However, the overall performance still confirms that viewpoint alignment plays a crucial role in improving operational effectiveness. Ablating all strategies led to a sharp decline, reducing the average success rate to 22%, underscoring the synergistic effect of combining these strategies for optimal performance.
We also evaluated our model with reduced demostrations. As shown in Table., reducing the number of demos from 30 to 15 resulted in a 21% drop in success rate. However, even with only 15 demos, the model slightly outperformed C2F with a single-camera setup trained on 30 demos, demonstrating its efficiency in leveraging limited demonstrations. Without any demos, the success rate dropped to 8%, highlighting the significant challenges of learning active vision-action skills in a large observation-action space. Reducing the number of demos also increased episode lengths and execution failures, emphasizing the importance of providing sufficient, high-quality samples early in training. Such samples can help the model quickly learn environmental geometry and facilitate more efficient action exploration in subsequent stages.
Interestingly, we observed that the model achieved meaningful ASSIG scores even without any demo data, slightly exceeding those obtained with 15 demos. We speculate that this is because when the global agent struggled to discover sparse task rewards, the local NBV agent may focused more on short-term entropy rewards $r_{\text{e}}$, leading to higher ASSIG scores. However, the AIIG scores, which require long-term planning, remained lower than those achieved with 15 demos.
@@
TABLE V: Ablation over different demo quantities.
Demo Result Proportion Interactivity Sensing
[…]
0 0.08 0.70 0.22 9.48 7.99 8.76 0.019 0.026
15 0.51 0.36 0.13 6.53 3.23 4.38 0.046 0.020
30 0.76 0.15 0.09 4.88 1.64 2.15 0.068 0.035